elimination / addition

-5x - 3y = -14
5x - 6y = 32
with steps too

Answer by mananth(16946)

-5x - 3y = -14
5x - 6y = 32
-5x-3y=-14 .............1
5x-6y=32 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
10x+6y=28
5x-6y=32
Add the two equations
15x=60
/15
x=4
plug value of x in (1)
-5x-3y=-14
-20-3y=-14
-3y=-14+ 20
-3y=6
y=-2